# Align-GRAG: Reasoning-Guided Dual Alignment for Graph Retrieval-Augmented Generation

### v2

@@ 5.11. Case study
In this case study, we investigate how summarization enhances the discovery of critical intermediate nodes within reasoning chains, as illustrated in Table reftab:summarization. Using well-crafted prompts in conjunction with a graph database, we show that concise, targeted summaries can surface pivotal entities—such as ‘J.K. Rowling’—that bridge the semantic gap between an initial query and the final answer. Our analysis evaluates the extent to which summarization reliably extracts these key connectors, thereby enabling coherent multi-hop reasoning that would otherwise be obscured within the raw graph structure. We further demonstrate that, without the interpretive layer provided by summarization, the underlying graph database lacks the necessary contextual signals to assemble the correct reasoning path end-to-end, underscoring the value of summarization as a catalyst for interpretable and effective knowledge retrieval.
5.12. LLM Prompts
@@
We introduce three prompts: the Summarization Prompt, the Generator Prompt, and the Relevance and Faithfulness Prompt. Each serves a specific function in enhancing the performance and evaluation of LLMs. Below is a detailed explanation of their purposes and applications, as illustrated in Figure 9, Figure 11, and Figure 10.
Summarization Prompt You are a helpful assistant responsible for generating a comprehensive summary of the data provided below. Given question and answer, and related graph data base. Please concatenate all of these into a single, comprehensive description. The description should logically connect the question to the answer. Make sure to include information collected from all descriptions. Question: {question} Answer: {Answer} Graph DataBase: {Textualized Graph}
Figure 9. Prompt for Summarization.
@@
The Summarization Prompt is specifically designed to generate a comprehensive and cohesive summary by synthesizing three key inputs: a given question, its corresponding answer, and the relevant textualized graph data. The goal of this prompt is to create a logical connection between the question and the answer, while also integrating information extracted from the provided graph data. This ensures the output not only answers the question but also demonstrates an understanding of the contextual relationships within the graph. By combining these elements, the prompt facilitates the generation of a rich, informative summary that serves as a bridge between structured and unstructured data.
@@
Relevance and Faithfulness Prompt Evaluation of Relevance: Evaluate the relevance of the Reasoning Chain in answering the QUESTION. A relevant Reasoning Chain contains information that helps answer the question, even if partially. Return one of the following labels: ’Relevant’, or ’Irrelevant’ without any additional response. Evaluation of Faithfulness: Evaluate the following Reasoning Chain for faithfulness in answering the QUESTION. A faithful response should include information that helps answer the question, even if partially, avoid inventing new details, and not contradict the context. Return one of the following labels: ’Faithful’ or ’Not Faithful’ without any additional response.
Figure 10. Evaluation of Relevance and Faithfulness Prompt.
@@
Figure 11. Prompt for Generator.
@@
  • Relevance Evaluation: This component assesses whether the reasoning chain provided is relevant to the given question. A relevant reasoning chain must contain information that contributes meaningfully to answering the question, even if the contribution is partial. The model is required to return one of two labels: Relevant or Irrelevant, with no additional commentary.
  • Faithfulness Evaluation: This component measures the faithfulness of the reasoning chain in responding to the question. A faithful reasoning chain should accurately represent the information in the context, avoid fabrications, and not contradict the provided data. Similarly, the model outputs one of two labels: Faithful or Not Faithful. This dual evaluation ensures the generated responses are both accurate and contextually grounded.
@@
6. Conclusion
This work introduces Align-GRAG, a novel framework designed to address the core challenges in graph-based retrieval-augmented generation: the inclusion of irrelevant knowledge during retrieval and the representation gap between graph embeddings and language model representations. To tackle these issues, Align-GRAG incorporates an innovative graph aligner module that performs dual alignment. The Graph Node Alignment mechanism prunes irrelevant nodes, preserving critical structural information, while Graph Representation Alignment bridges the semantic gap, ensuring a unified representation space between graph and language models. Extensive experiments on the GraphQA benchmark demonstrate that Align-GRAG consistently outperforms existing baselines across multiple tasks, such as commonsense reasoning and knowledge graph understanding. By effectively integrating graph reasoning with LLM-based generation, Align-GRAG achieves more accurate, context-aware, and reliable outputs. This work paves the way for future advancements in GRAG systems, offering a robust solution for leveraging structured graph data in retrieval-augmented generation tasks.
7. Limitations
Despite the promising results demonstrated by Align-GRAG, this work has certain limitations. Due to resource constraints, we were unable to conduct experiments on larger LLMs, leaving the effectiveness of the proposed alignment approach on more powerful models uncertain. Additionally, since our method requires the generation and utilization of graph embeddings, it cannot be directly implemented on closed-source models, such as GPT-4, which restricts access to internal embedding representations. These limitations highlight potential areas for future exploration, such as validating the scalability of the approach with state-of-the-art LLMs and developing techniques to adapt Align-GRAG for closed-source environments.

### v3

@@ Appendix D Case study
In this case study, we investigate how summarization enhances the discovery of critical intermediate nodes within anchors and rationales, as illustrated in Table reftab:summarization. Using well-crafted prompts in conjunction with a graph database, we show that concise, targeted summaries can surface pivotal entities—such as ‘J.K. Rowling’—that bridge the semantic gap between an initial query and the final answer. Our analysis evaluates the extent to which summarization reliably extracts these key connectors, thereby enabling coherent multi-hop reasoning that would otherwise be obscured within the raw graph structure. We further demonstrate that, without the interpretive layer provided by summarization, the underlying graph database lacks the necessary contextual signals to assemble the correct reasoning path end-to-end, underscoring the value of summarization as a catalyst for interpretable and effective knowledge retrieval.
Appendix E LLM Prompts
@@
We introduce three prompts: the Extraction Prompt, the Generator Prompt, and the Relevance and Faithfulness Prompt. Each serves a specific function in enhancing the performance and evaluation of LLMs. Below is a detailed explanation of their purposes and applications, as illustrated in Figure 10, Figure 12, and Figure 11.
Anchor and Rationale Extraction Prompt You are a helpful assistant for anchor and rationale extraction from graph evidence. Inputs: a question, its answer, and a textualized graph database (nodes/edges described in text). Goal: extract (1) a rationale chain that links the question to the answer using graph evidence, and (2) a small set of anchors (key intermediate entities/relations) that can be grounded to graph nodes/edges. Definitions: Rationale chain: an ordered list of short reasoning steps. Each step should mention the evidence used and move closer to the answer. Anchors: reasoning-critical intermediate entities/relations that appear in the graph database and are necessary for the reasoning. Constraints: Use only information supported by the provided graph database and the given answer. Anchors must be verbatim spans from the graph database (copy exact surface forms). Keep each rationale step concise. Output format (follow exactly): 1. RationaleChain: a numbered list of steps (3–6 steps). 2. Anchors: a bullet list of anchors. For each anchor, provide its type (entity or relation) and the copied span. Question: {question} Answer: {Answer} Graph DataBase: {Textualized Graph} Now produce the output.
Figure 10: Prompt for anchor and rationale extraction.
@@
The Extraction Prompt is specifically designed to generate a comprehensive and cohesive anchor and rationale by synthesizing three key inputs: a given question, its corresponding answer, and the relevant textualized graph data. The goal of this prompt is to create a logical connection between the question and the answer, while also integrating information extracted from the provided graph data. This ensures the output not only answers the question but also demonstrates an understanding of the contextual relationships within the graph. By combining these elements, the prompt facilitates the generation of a rich, informative summary that serves as a bridge between structured and unstructured data.
@@
Relevance and Faithfulness Prompt Evaluation of Relevance: Evaluate the relevance of the anchor and rationale in answering the QUESTION. The relevant anchor and rationale contain information that helps answer the question, even if partially. Return one of the following labels: ’Relevant’, or ’Irrelevant’ without any additional response. Evaluation of Faithfulness: Evaluate the following anchor and rationale for faithfulness in answering the QUESTION. A faithful response should include information that helps answer the question, even if partially, avoid inventing new details, and not contradict the context. Return one of the following labels: ’Faithful’ or ’Not Faithful’ without any additional response.
Figure 11: Evaluation of Relevance and Faithfulness Prompt.
@@
Figure 12: Prompt for Generator.
@@
  • Relevance Evaluation: This component assesses whether the anchor and rationale provided is relevant to the given question. A relevant anchor and rationale must contain information that contributes meaningfully to answering the question, even if the contribution is partial. The model is required to return one of two labels: Relevant or Irrelevant, with no additional commentary.
  • Faithfulness Evaluation: This component measures the faithfulness of anchor and rationale in responding to the question, which should accurately represent the information in the context, avoid fabrications, and not contradict the provided data. Similarly, the model outputs one of two labels: Faithful or Not Faithful. This dual evaluation ensures the generated responses are both accurate and contextually grounded.